OpenStudy (anonymous):

hello everyone can some one help me out I have to creat a Create a linear equation which requires one step to solve

OpenStudy (saifoo.khan):

Sure, where's the question?

OpenStudy (anonymous):

I have to make a linear equTION WHICH requires one step to solve I have to make my own.

OpenStudy (saifoo.khan):

Ah! Okay.. So, you want a question then solve the answer to make: y = mx + b

OpenStudy (anonymous):

yes and I also need one that you can solve in two stpes to

OpenStudy (saifoo.khan):

Ok, so let's begin. Let's say u have two points and u want to find the equation. Points = (0,1) and (2,5) Is this question fine?

OpenStudy (anonymous):

um Im not for sure dont know how to do this at all

OpenStudy (anonymous):

I can give you an exsample if that would help ?

OpenStudy (saifoo.khan):

This is also a question i just posted. if u want your own that will be fine as well.

OpenStudy (anonymous):

2x = 10 here is the exsample

OpenStudy (saifoo.khan):

Ok, so first of all we will solve it, 2x = 10 x = 10/2 x = 5 Now we have x = 5, so we will graph it now,